OpenStudy (anonymous):

Find the product: 2 1/8 × 1 1/6 × 3

OpenStudy (anonymous):

PLEASE HELP!!

OpenStudy (anonymous):

Convert to improper fractions. Multiply numerators and denominators.

OpenStudy (anonymous):

2 1/8 = 17/8 1 1/6 = 7/6

OpenStudy (anonymous):

\[\left(2\frac{1}{8}\right)*\left(1\frac{1}{6}\right)*3=\frac{119}{16}=7\frac{7}{16} \]